OpenStudy (angelax0):

The length of a rectangle is 5 m less than twice the width, and the area of the rectangle is 52 m^2 . Find the dimensions of the rectangle.

OpenStudy (jackthegreatest):

(2w-5)w=52

OpenStudy (jackthegreatest):

2w^2-5w=52

OpenStudy (jackthegreatest):

2w^2-5w-52=0

OpenStudy (jackthegreatest):

Factor... (2w-13)(w+4)

OpenStudy (jackthegreatest):

2w=13 or w = -4

OpenStudy (jackthegreatest):

Width can't be negative

OpenStudy (jackthegreatest):

W=6.5 m

OpenStudy (jackthegreatest):

@angelax0